After taking off from the landing pad, a helicopter flies 9 miles south, then 8 miles northeast. How far is the helicopter from the landing pad (s), and in what direction (pheta) must it fly to head directly back to the landing pad?

s= and
pheta = (0 degree <= pheta < 360 where East is 0 degree and North is 90 degree).
1
Expert's answer
2019-07-30T10:01:42-0400

Using the cosine rule,

x2=82+92-2*8*9cos 450

x2=69.35

x=8.328 miles from s

9/sin y=8.328/sin450

sin y=9sin450/8.328=0.9196

y=arcsin 0.9196=66.870.

To get back to s it travels in the direction 1800+(66.870-450)=201.870.